OpenStudy (anonymous):

need help on number 11 is itright or wrong if wrong shom me how to work it out

OpenStudy (anonymous):

second page

OpenStudy (anonymous):

compute \[\frac{1}{4}\times 800+\frac{3}{4}\times 200\]

OpenStudy (anonymous):

i get 350 which is less than the sure thing of $400

OpenStudy (anonymous):

so expected value of M is LESS THAN expected value of N

OpenStudy (anonymous):

satellite did u see "Be_creepy" anywhere? or ivy or persephone? today?

OpenStudy (anonymous):

soo its A

OpenStudy (anonymous):

yeah, A

OpenStudy (anonymous):

yessss!!!!!